OpenStudy (yacoub1993):

Complete the square and write the equation in standard form. Then give the center and radius of the circle. x2 + y2 - 10x - 8y + 29 = 0

myininaya (myininaya):

\[(x^2-10x)+(y^2-8y)=-29 \] \[(x^2-10x+?_1)+(y^2-8y+?_2)=-29+?_1+?_2 \] We need to figure out what those ?'s should be

myininaya (myininaya):

To complete the square we do \[x^2 \pm kx+ (\frac{k}{2})^2=(x \pm \frac{k}{2})^2 \]

myininaya (myininaya):

\[y^2\pm ky+(\frac{k}{2})^2=(y \pm \frac{k}{2})^2\]

myininaya (myininaya):

So take the number in front of x and then divide it by 2 then square that result that will be ?_1

OpenStudy (yacoub1993):

Q. what does k stand for cause in the equation i gave there is no k

myininaya (myininaya):

\[x^2-10x+?_1 \] \[x^2-kx+(\frac{k}{2})^2\]

myininaya (myininaya):

do you see what k is now?

myininaya (myininaya):

I'm giving you a formula for completing the square I'm trying to get you to identify what k is in the expression I just wrote.

myininaya (myininaya):

What is in front of x in both expressions?

myininaya (myininaya):